# Learning relevant contextual variables within Bayesian optimization

### v3

@@ 5.3 Sensitivity analysis
We now report experiments assessing the robustness of SADCBO’s performance to several modifications, either at the hyperparameter level or at the experiment setting level. The latter includes assessing performance when increasing the number of noise variables, varying the contextual variable query cost, or varying the surrogate model.
@@
We investigate four different values for the query cost of contextual variables (Figure S4). For extremely cheap contextual variables $\lambda_{j}=0.1$ for all $j$, that is, ten times cheaper than a design variable, VBO performs favorably, as optimizing over all inputs $[\mathbf{x},\mathbf{z}]$ is cheap. SADCBO remains competitive in this configuration, even though MMDBO and CaBO can handle this case. For a moderate cost $\lambda_{j}=1$ (the cost model considered in Figure 2), SADCBO obtains the lowest average rank over all four test functions. For expensive contextual variables, $\lambda_{j}=3$ or $\lambda_{j}=10$, CaBO seems overall more suitable, although closely followed by SADCBO, and CBO.
@@
As SADCBO relies on a posterior sensitivity analysis to select the relevant contextual variables, and is hence agnostic to the choice of GP surrogate model, it can be combined with other methods that induce sparsity via the GP surrogate. One such method is by Eriksson and Jankowiak (2021), who introduced a sparsity-enforcing GP surrogate equipped with a horseshoe prior on the square inverse lengthscales, coined SAASBO. In Figure 3, we compare SADCBO with the combined method SAASBO+SADCBO, with both having the same hyperparameters. We observe that SAASBO+SADCBO improves over just SAASBO in all the synthetic examples, and is also better than SADCBO in two out of four examples. Note that the performance of SAASBO+SADCBO may further improve through hyperparameter tuning.
Figure 3: Combining SADCBO with sparsity-enforcing surrogate SAASBO. For any variable, the associated query cost is 1. $p(\mathbf{z})=\mathcal{U}([0,1]^{c})$. The combination is fruitful and improves the performances of SAASBO (green vs blue curves).
@@
We vary the 3 hyperparameters of SADCBO: $\eta,\gamma,Q$. Unsurprisingly, the cumulative sensitivity threshold $\eta$ stands out as the most relevant parameter: as its value decreases, fewer variables are included, at which point not all relevant ones are selected, leading to reduced performance (Appendix C).
@@ 6 Conclusion
In this paper, we extended Contextual BO (Krause and Ong, 2011) to settings in which the contextual variables may be not only observed but also optimized at a cost. We introduced SADCBO, an algorithm designed to select relevant context variables affecting the experimental outcomes by efficiently leveraging information present in both the observational and the interventional data. SADCBO results in more adequate surrogate models, and ensures the reproducibility of experiments by controlling for such relevant variables. In that respect, SADCBO should be used for practical applications where contextual variables can have an influence while being controllable. This includes, e.g., the development of new high-throughput materials or drugs, where machine learning strategies are being increasingly used (Zhang et al., 2020; Gómez-Bombarelli et al., 2018). SADCBO can also be combined with any GP surrogate. Thus, if a practitioner believes that a specific contextual variable should be included, this can be easily achieved. Conversely, the variable selection procedure could be generalized to discard design variables as well.
@@
To achieve cost efficiency, SADCBO integrates the query cost at the variable selection level and employs an early stopping criterion. The latter only depends on an upper bound on the instantaneous regret difference and is therefore not cost-aware. Adding a notion of remaining budget to this criterion would certainly benefit our approach. On a similar note, while our algorithm allows incorporating cost, more effort could be put into specifying the costs. In our experiments, they were set to $1$ for all variables to prevent bias in the results, and we carried out an ablation study with different costs in Section 5.3. An interesting avenue for future work would be to elicit knowledge of experimental costs from domain experts in real-world situations.
Supplementary Material
Learning relevant contextual variables within Bayesian optimization
@@
The Appendix is organized as follows. In Appendix A, we provide further details about the Stopping Criterion introduced in Section 3.2. In Appendix B, we provide more details about one of the baselines used in the main text, based on maximum mean discrepancy. Appendix C contains further experimental results regarding:
  • The distribution of early stopping time for SADCBO (Figure S1).
  • Stopping time criterion and sensitivity-based inclusion probabilities of contextual variables found by SADCBO for additional test functions (Figure S2).
@@
  • Varying SADCBO hyperparameters (Section C.1).
@@
Finally, Appendix D contains a description of the real-world experiments performed throughout the paper, along with the analytical expressions of the synthetic examples used.
Appendix A Stopping Criterion
The stopping criterion we use is derived from Ishibashi et al. (2023). The absolute difference of expected minimum simple regrets $\Delta R_{t}:=\lvert R_{t}-R_{t-1}\rvert$ can be upper bounded with probability $1-\delta$ by $\Delta\tilde{R}_{t}$, a quantity defined in Eq. 10, which involves a sequence $\kappa_{\delta,t-1}$:
$\kappa_{\delta,t-1}=\max_{\mathbf{v}\in\mathcal{D}_{t-1}}~{}\text{UCB}_{\delta}(\mathbf{v})-\max_{\mathbf{v}\in\mathcal{V}}~{}\text{LCB}_{\delta}(\mathbf{v}),$ (S1)
@@ Appendix A Stopping Criterion
where $\text{UCB}_{\delta}(\mathbf{v})=\mu_{t}(\mathbf{v}|\mathcal{D}_{t})+\beta_{t}^{1/2}\sigma_{t}(\mathbf{v}|\mathcal{D}_{t})$ and $\text{LCB}_{\delta}(\mathbf{v})=\mu_{t}(\mathbf{v}|\mathcal{D}_{t})-\beta_{t}^{1/2}\sigma_{t}(\mathbf{v}|\mathcal{D}_{t})$. $\beta_{t}^{1/2}$ is a trade-off parameter between exploration and exploitation that depends on $\delta$ (Srinivas et al., 2012). $\kappa_{\delta,t-1}$ is a quantity first introduced by Makarova et al. (2022, Section 3.2) as an upper bound for the simple regret of the surrogate, which directly flows from the bounds provided by Srinivas et al. (2012) for well-calibrated surrogates.
Next, again following Ishibashi et al. (2023), The upper bound $\Delta\tilde{R}_{t}$ can itself be upper bounded by $\Delta\hat{R}_{t}$, a quantity whose convergence speed is limited by $s_{t}$ (Equation 11), hence why the stopping criterion $\Delta\tilde{R}_{t}\leq s_{t}$.
Appendix B Maximum Mean Discrepancy-based variable selection
@@
Spagnol et al. (2019) introduced a BO algorithm with a variable selection procedure based on the Hilbert Schmidt Independence Criterion (HSIC). This measure can be used in our setting as well. We now briefly describe how it is defined.
@@
As introduced in the main text, let $\mathcal{Z}\subset\mathbb{R}^{c}$ be the space of contextual variables, and $\mathcal{H}$ be a Hilbert space of $\mathbb{R}$-valued functions on $\mathcal{Z}$. Assume that $k:\mathcal{Z}\times\mathcal{Z}\to\mathbb{R}$ is the unique positive definite kernel associated with the Reproducing Kernel Hilbert Space $\mathcal{H}$. Let $\mu_{\mathbb{P}_{Z}}$ be the kernel mean embedding of the distribution $\mathbb{P}_{Z}$, $\mu_{\mathbb{P}_{Z}}:=\mathbb{E}_{Z}[k(Z,\cdot)]=\int_{\mathcal{Z}}k(\mathbf{z},\cdot)\mathrm{d}\mathbb{P}_{Z}$. Kernel embeddings of probability measures provide a distance between distributions between their embeddings in the Hilbert Space $\mathcal{H}$, named Maximum Mean Discrepancy (MMD, (Gretton et al., 2012)):
@@
$~{}~{}~{}+\mathbb{E}_{Z}\mathbb{E}_{Y}\mathbb{E}_{Z^{\prime}}\mathbb{E}_{Y^{\prime}}[k(Z,Z^{\prime})l(Y,Y^{\prime})]$
$~{}~{}~{}-2\mathbb{E}_{Z,Y}\mathbb{E}_{Z^{\prime}}\mathbb{E}_{Y^{\prime}}[k(Z,Z^{\prime})l(Y,Y^{\prime})].$
To determine the relevance of a variable $Z^{(i)}$, Spagnol et al. (2019) introduce
@@
Appendix C Additional experimental results
Figure S1: Distribution of stopping criterion triggering time for SADCBO across $N=100$ different BO trials. We consider the Ackley5D function with an increasingly larger ratio of relevant contextual variables over design variables, and 8 irrelevant contextual variables. $p(\mathbf{z})=\mathcal{U}([0,1]^{c})$. For any variable, the associated query cost is 1. As the impact of contextual variables on the output function grows, the number of iterations spent in the observational phase grows as well.
Figure S2: Each row deals with a specific problem. The left panel shows the BO trial results for each baseline. The middle and right panel show statistics related to SADCBO: 1) the stopping time at which phase II starts and 2) the inclusion probabilities for each contextual variable. Statistics are computed across $N=100$ BO trials with different seeds.
@@ Appendix C Additional experimental results
C.1 Additional details on hyperparameter variations.
@@
Appendix D Experiment details
D.1 Real-world datasets
@@
This dataset was first introduced in (Cakmak et al., 2020). The goal is to tune the hyper-parameters of a trading strategy so as to maximize return under risk-aversion to random environmental conditions. A software is used to simulate and optimize the evolution of a portfolio over a period of four years using open-source market data. Each evaluation of this simulator returns the average daily return over this period of time under the given combination of hyper-parameters and environmental conditions. Since the simulator is expensive to evaluate, we do not use it directly but perform pool-based Bayesian Optimization using a pool of 3000 points generated according to a Sobol sampling design.
@@
This dataset comes from the UCI Machine Learning Repository (Gerritsma et al., 2013). The optimization problem is to maximize the residuary resistance per unit weight of displacement of a yacht by controlling its 5-dimensional hull geometry coefficients. Another optimization variable is the 1-dimensional Froude number. We chose as design variables the first four dimensions of the hull geometry coefficients. The contextual variables are the last hull geometry dimension and the Froude number. Like the Portfolio optimization dataset, we have access to a limited number of samples $(\approx 300)$ and thus perform pool-based Bayesian optimization.
@@
[…]
@@
This task was first introduced in Wang et al. (2017), and consists of a control parameter tuning problem for robot pushing. This real-world function returns the distance between a designated goal location and two objects being pushed by two robot hands, whose trajectory is determined by 14 parameters specifying the location, rotation, velocity and moving direction, among others. The function is implemented with a physics engine, the Box2D simulator. There are 6 design variables and 8 contextual variables.
D.2 Synthetic test functions

### v4

@@ 5.3 Sensitivity analysis
We now report experiments assessing the robustness of SADCBO’s performance to several modifications, either at the hyperparameter level or at the experiment setting level. The latter includes assessing performance when increasing the number of noise variables, varying the contextual variable query cost, or varying the surrogate model. Next, additional experiments illustrate the sound behavior of the proposed phase switching criterion implemented in SADCBO.
@@
We investigate four different values for the query cost of contextual variables (Figure S4). For extremely cheap contextual variables $\lambda_{j}=0.1$ for all $j$, that is, ten times cheaper than a design variable, VBO performs favorably, as optimizing over all inputs $[\mathbf{x},\mathbf{z}]$ is cheap. SADCBO remains competitive in this configuration, even though MMDBO and CaBO perform on par. For a moderate cost $\lambda_{j}=1$ (the cost model considered in Figure 2), SADCBO obtains the lowest average rank over all four test functions. For expensive contextual variables, $\lambda_{j}=3$ or $\lambda_{j}=10$, CaBO seems overall more suitable, although closely followed by SADCBO, and CBO.
@@
As SADCBO relies on a posterior sensitivity analysis to select the relevant contextual variables, and is hence agnostic to the choice of GP surrogate model, it can be combined with other methods that induce sparsity via the GP surrogate. One such method is by Eriksson and Jankowiak [2021], who introduced a sparsity-enforcing GP surrogate equipped with a horseshoe prior on the square inverse lengthscales, coined SAASBO. In Figure 3, we compare SADCBO with the combined method SAASBO+SADCBO, with both having the same hyperparameters. We observe that SAASBO+SADCBO improves over just SAASBO in all the synthetic examples, and is also better than SADCBO in two out of four examples. Note that the performance of SAASBO+SADCBO may further improve through hyperparameter tuning.
Figure 3: Combining SADCBO with sparsity-enforcing surrogate SAASBO. For any variable, the associated query cost is 1. $p(\mathbf{z})=\mathcal{U}([0,1]^{c})$. The combination is fruitful and improves the performances of SAASBO.
Figure 4: Assessing SADCBO’s phase switching criterion on the Hartmann6D function. The iteration selected by the adaptive stopping criterion implemented in SADCBO yields one of the best BO trials. Each curve is computed as an average of 10 different random seeds.
SADCBO phase switching criterion.
We ensure that the criterion is well-behaved: the more information about the output is contained in the contextual variables, the later the phase switching occurs (Figure S5). Even though the stopping criterion was initially devised for vanilla BO, its application in a CBO setting is fruitful. Figure 4 further illustrates the soundness of the phase switching criterion. Using the Hartmann6D function under the same setting as described above, the mean switching iteration found by SADCBO over 100 different runs was collected. Then, new BO runs using SADCBO with a fixed phase switching time $i\in\{1,\dots,100\}$ were performed. This was done 10 times for each switching time, using different random seeds for the initial dataset. The switching time found by SADCBO yields one of the best runs, validating the use of the criterion.
@@
We vary the 3 hyperparameters of SADCBO: $\eta,\gamma,Q$. Unsurprisingly, the cumulative sensitivity threshold $\eta$ stands out as the most relevant parameter: as its value decreases, fewer variables are included, at which point not all relevant ones are selected, leading to reduced performance (Appendix D).
@@ 6 CONCLUSION
In this paper, we extended Contextual BO [Krause and Ong, 2011] to settings in which the contextual variables may be not only observed but also optimized at a cost. We introduced SADCBO, an algorithm designed to select relevant context variables affecting the experimental outcomes by efficiently leveraging information present in both the observational and the interventional data. SADCBO results in more adequate surrogate models, and ensures the reproducibility of experiments by controlling for such relevant variables. In that respect, SADCBO should be used for practical applications where contextual variables can have an influence while being controllable. This includes, e.g., the development of new high-throughput materials or drugs, where machine learning strategies are being increasingly used [Zhang et al., 2020, Gómez-Bombarelli et al., 2018]. SADCBO can also be combined with any GP surrogate. Thus, if a practitioner believes that a specific contextual variable should be included, this can be easily achieved. Conversely, the variable selection procedure could be generalized to discard design variables as well. Lastly, recent work [Branchini et al., 2023] proposed to perform BO under the assumption that the input variables and the output are linked by a causal directed acyclic graph, learning the graph whilst maximizing the objective function. Despite its high computational complexity, applying this technique to our particular problem might be promising.
@@
To achieve cost efficiency, SADCBO integrates the query cost at the variable selection level and employs an early stopping criterion. The latter only depends on an upper bound on the instantaneous regret difference and is therefore not cost-aware. Adding a notion of remaining budget to this criterion would certainly benefit our approach. On a similar note, while our algorithm incorporates cost, more effort could be put into specifying the costs. In our experiments, they were set to $1$ for all variables to prevent bias in the results, and we carried out an ablation study with different costs in Section 5.3. Yet, it is worth mentioning that our method is compatible with the inference of black-box, input-dependent costs, similarly to CaBO [Lee et al., 2020]. One would simply need to modify Equation 9, replacing $\lambda_{j}$ by the learned cost. An interesting avenue for future work would be to elicit knowledge of experimental costs from domain experts in real-world situations.
Acknowledgements.
JM acknowledges the support of the Research Council of Finland under the HEALED project (grant 13342077). AB, AT, STJ, and PR were supported by the Research Council of Finland Flagship programme: Finnish Center for Artificial Intelligence FCAI. AT further acknowledges funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No. 101059891. SJS and SK were supported by the UKRI Turing AI World-Leading Researcher Fellowship, [EP/W002973/1].
Appendix
@@ Acknowledgements.
The Appendix is organized as follows. In Appendix A, we provide a flowchart summarizing the proposed method SADCBO. In Appendix B, we provide further details about the phase switching criterion introduced in Section 3.2. In Appendix C, we provide more details about one of the baselines used in the main text, based on maximum mean discrepancy. Appendix D contains further experimental results regarding:
  • Phase switching time and sensitivity-based inclusion probabilities of contextual variables found by SADCBO for additional test functions (Figure S2).
@@
  • The distribution of phase switching times for SADCBO (Figure S5).
  • Varying SADCBO hyperparameters (Section D.1 and Figure S6).
@@
Finally, Appendix E contains a description of the real-world experiments performed throughout the paper, along with the analytical expressions of the synthetic examples used.
Appendix A Flowchart of the algorithm
Figure S1: Flowchart of the proposed method SADCBO.
Appendix B Phase Switching Criterion
The phase switching criterion we employ is derived from the stopping criterion from Ishibashi et al. [2023]. The absolute difference of expected minimum simple regrets $\Delta R_{t}:=\lvert R_{t}-R_{t-1}\rvert$ can be upper bounded with probability $1-\delta$ by $\Delta\tilde{R}_{t}$, a quantity defined in Equation 11. Directly quoting the work of Ishibashi et al. [2023], the rationale behind this criterion reads as follows: “By evaluating the difference between the expected minimum simple regrets, we can stop BO without knowing $f^{*}$ , because it indicates that the search efficiency is low and there is almost no improvement in the objective value. However, it is generally difficult to calculate $\Delta R_{t}$ analytically”. Next, any stopping criterion involves the computation of some sort of threshold. Ishibashi et al. [2023] exploit the fact that their upper bound $\Delta\tilde{R}_{t}$ can itself be upper bounded by a quantity (introduced in [Ishibashi et al., 2023, Equation 10]), whose convergence speed to zero is limited by a specific term, $s_{t}$ (Equation 12). $s_{t}$ can be computed analytically and therefore yields an adaptive threshold.
Finally, Equation 11 involves a sequence $\kappa_{\delta,t-1}$:
$\kappa_{\delta,t-1}=\max_{\mathbf{v}\in\mathcal{D}_{t-1}}\leavevmode\nobreak\ \text{UCB}_{\delta}(\mathbf{v})-\max_{\mathbf{v}\in\mathcal{V}}\leavevmode\nobreak\ \text{LCB}_{\delta}(\mathbf{v}),$ (S1)
@@ Appendix B Phase Switching Criterion
where $\text{UCB}_{\delta}(\mathbf{v})=\mu_{t}(\mathbf{v}|\mathcal{D}_{t})+\beta_{t}^{1/2}\sigma_{t}(\mathbf{v}|\mathcal{D}_{t})$ and $\text{LCB}_{\delta}(\mathbf{v})=\mu_{t}(\mathbf{v}|\mathcal{D}_{t})-\beta_{t}^{1/2}\sigma_{t}(\mathbf{v}|\mathcal{D}_{t})$. $\beta_{t}^{1/2}$ is a trade-off parameter between exploration and exploitation that depends on $\delta$ [Srinivas et al., 2012]. $\kappa_{\delta,t-1}$ is a quantity that was first introduced by Makarova et al. [2022, Section 3.2] as an upper bound for the simple regret of the surrogate, which directly flows from the bounds provided by Srinivas et al. [2012] for well-calibrated surrogates.
Heuristically, one can think of our setting as applying the stopping criterion to $\mathbf{x}\mapsto f(\mathbf{x},\mathbf{z})$, a stochastic black-box function with $\mathbf{z}\sim p(\mathbf{z})$. Upon satisfaction of this criterion, we switch to the optimization of $(\mathbf{x},\mathbf{z})\mapsto f(\mathbf{x},\mathbf{z})$ where some contextual variables are optimized, and some others are still sampled from $p(z^{(j)})$.
Appendix C Maximum Mean Discrepancy-based variable selection
@@
Spagnol et al. [2019] introduced a BO algorithm with a variable selection procedure based on the Hilbert Schmidt Independence Criterion (HSIC). This measure can be used in our setting as well. We now briefly describe how it is defined.
@@
As introduced in the main text, let $\mathcal{Z}\subset\mathbb{R}^{c}$ be the space of contextual variables, and $\mathcal{H}$ be a Hilbert space of $\mathbb{R}$-valued functions on $\mathcal{Z}$. Assume that $k:\mathcal{Z}\times\mathcal{Z}\to\mathbb{R}$ is the unique positive definite kernel associated with the Reproducing Kernel Hilbert Space $\mathcal{H}$. Let $\mu_{\mathbb{P}_{Z}}$ be the kernel mean embedding of the distribution $\mathbb{P}_{Z}$, $\mu_{\mathbb{P}_{Z}}:=\mathbb{E}_{Z}[k(Z,\cdot)]=\int_{\mathcal{Z}}k(\mathbf{z},\cdot)\mathrm{d}\mathbb{P}_{Z}$. Kernel embeddings of probability measures provide a distance between distributions between their embeddings in the Hilbert Space $\mathcal{H}$, named Maximum Mean Discrepancy (MMD, [Gretton et al., 2012]):
@@
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ +\mathbb{E}_{Z}\mathbb{E}_{Y}\mathbb{E}_{Z^{\prime}}\mathbb{E}_{Y^{\prime}}[k(Z,Z^{\prime})l(Y,Y^{\prime})]$
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ -2\mathbb{E}_{Z,Y}\mathbb{E}_{Z^{\prime}}\mathbb{E}_{Y^{\prime}}[k(Z,Z^{\prime})l(Y,Y^{\prime})].$
To determine the relevance of a variable $Z^{(i)}$, Spagnol et al. [2019] introduce
@@
Appendix D Additional experimental results
Figure S2: Each row deals with a specific problem. The left panel shows the BO trial results for each baseline. The middle and right panel show statistics related to SADCBO: 1) the phase switching time after which phase II begins and 2) the inclusion probabilities for each contextual variable. Statistics are computed across $N=100$ BO trials with different seeds.
@@ Appendix D Additional experimental results
Figure S5: Distribution of phase switching criterion triggering times for SADCBO across $N=100$ different BO trials. We consider the Ackley5D function with an increasingly larger ratio of relevant contextual variables over design variables, and 8 irrelevant contextual variables. $p(\mathbf{z})=\mathcal{U}([0,1]^{c})$. For any variable, the associated query cost is 1. As the impact of contextual variables on the output function grows, the number of iterations spent in the observational phase grows as well.
D.1 Additional details on hyperparameter variations.
@@
Appendix E Experiment details
E.1 Real-world datasets
@@
This dataset was first introduced in [Cakmak et al., 2020]. The goal is to tune the hyper-parameters of a trading strategy so as to maximize return under risk-aversion to random environmental conditions. A software is used to simulate and optimize the evolution of a portfolio over a period of four years using open-source market data. Each evaluation of this simulator returns the average daily return over this period of time under the given combination of hyper-parameters and environmental conditions. Since the simulator is expensive to evaluate, we do not use it directly but perform pool-based Bayesian Optimization using a pool of 3000 points generated according to a Sobol sampling design.
@@
This dataset comes from the UCI Machine Learning Repository [Gerritsma et al., 2013]. The optimization problem is to maximize the residuary resistance per unit weight of displacement of a yacht by controlling its 5-dimensional hull geometry coefficients. Another optimization variable is the 1-dimensional Froude number. We chose as design variables the first four dimensions of the hull geometry coefficients. The contextual variables are the last hull geometry dimension and the Froude number. Like the Portfolio optimization dataset, we have access to a limited number of samples $(\approx 300)$ and thus perform pool-based Bayesian optimization.
@@
This case is a computational chemistry challenge. Molecules can adopt different structures that preserve the topology (bonds and bonding types), but have different internal angles. Finding such conformers is a global optimization problem. Here, we are searching for the conformers of alanine — a molecule with structure C3H7NO2 — whose energy is calculated at each round of BO with the AMBER force field [Salomon Ferrer et al., 2013, Case et al., 2023]. Alanine provides 33 structural variables to optimize: ten dihedral angles, eleven bond angles, and twelve bond lengths. Conformer search in the full 33-dimensional space is very challenging, but progress has been made with Bayesian optimization recently by reducing the problem to the four most important dihedral angles [Fang et al., 2021]. For the example in this work, three of these four dihedral angles were chosen as the design variables (indices 3, 17, and 21 in the dataset; which denotes dihedral angles d4, d11, and d13 in AMBER notation; d4 is the bond leading to the amino group, d13 the one leading to the hydroxyl group, and d11 is the bond between these two), the rest of the dihedral and bond angles (18 angles) are chosen as the contextual variables, and the bond lengths are kept fixed to facilitate faster simulations. The search space is selected by utilizing molecule domain knowledge in a conservative manner that allows 10-20 degree variations for the bond angles and is free for the dihedral angles. To outline the alanine optimization results, the structure optimization performed here as a test case is a high-dimensional problem, thus, the VBO method that tries to optimize all the variables $\mathbf{x}$ and $\mathbf{z}$ converges slowly. Due to the same reason, methods MMDBO and Dropout also underperform in terms of convergence for the alanine problem. Similarly to SADCBO, these two baselines operate variable selection, although using a different selection criterion. However, controlling the selected variable comes at a cost. On the opposite, it turns out that from Figure S2 (fifth row, middle panel), SADCBO virtually never switches to phase II for the Molecular Shape example. Therefore, SADCBO does perform contextual variable selection, but does not control them, it only chooses which of these variables will be included in the surrogate, hence behaving like CBO, but with a variable selection step. This explains why 1) CBO closely follows SADCBO for this example and 2) why other variable selection baselines like MMDBO and Dropout end up far from SADCBO. Interestingly, in this case, the simplified case of optimizing only the design variables $\mathbf{x}$ (CUBO) also performs well. This is because our domain experts made good initial choices on the relevant design variables $\mathbf{x}$ and the search spaces of context variables $\mathbf{z}$. This type of pre-analysis is time-consuming and more challenging for larger molecules. Hence, a future line of work is to test context-aware BO more comprehensively in molecule structure optimization tasks.
@@
This task was first introduced in Wang et al. [2017], and consists of a control parameter tuning problem for robot pushing. This real-world function returns the distance between a designated goal location and two objects being pushed by two robot hands, whose trajectory is determined by 14 parameters specifying the location, rotation, velocity and moving direction, among others. The function is implemented with a physics engine, the Box2D simulator. There are 6 design variables and 8 contextual variables.
E.2 Synthetic test functions